### v1

@@ 3.2 FT-Mamba
Following the design of FT-Transformer (Gorishniy et al., 2021), we adapt the Mamba architecture (Gu & Dao, 2023) for tabular feature extraction and name it as FT-Mamba (Feature Tokenizer + Mamba). As shown in Fig. 1 (a), we first separately tokenize the categorical and numerical tabular features to embeddings and add a [CLS] token at the sequence end. Then a stack of mamba blocks is used to extract the high-dimensional features from the embeddings, and the final representation of the [CLS] token is used as the original tabular embedding, $\boldsymbol{z}_{tab}$.
@@ 3.2.1 Feature Tokenizer
Our feature tokenizer module follows the design of (Gorishniy et al., 2021). It transfers the tabular input $\boldsymbol{x}_{tab}$ to embeddings $\boldsymbol{t}_{tab}\in\mathbb{R}^{N\times D_{tab}}$ as shown in Fig. 2 (a). For continuous features, the feature value is mapped by a linear function with $D_{tab}$-dimensional weights and bias. For categorical features, each category is mapped to a vector through a lookup table and is added with the corresponding bias.
@@ 3.3.1 Problem Formulation
To address the cross-modal discrepancies in feature dimensionality and confidence levels, we first quantify the characteristics of unimodal features. Since the features have been processed by the deep encoders, redundant or homogeneous information with low entropy has already been discarded. Therefore, we use the $\ell_{1}$-norm, which intuitively reflects the sparsity of feature vectors, to measure the information content of a unimodal feature as in (1). It reflects the feature’s informational contribution to the classification decision.
@@
Then we the define feature density of a feature vector as the ratio of its content to feature dimension, as shown in (2). It describes the average information content on each element and reflects the overall amplitude. Equation (1) and (2) use $\boldsymbol{z}_{img}$ as an example, and the same concept applies to the tabular modality.
@@
Additionally, we introduce the modality confidence ratio $R_{\text{conf}}$ to guide the reliability of each modality in downstream tasks. This ratio is defined as the confidence level of image features relative to tabular features in the current downstream task, as shown in (3). In our experiments, $R_{\text{conf}}$ is calculated by the unimodal performance of pretrained encoders.
[…]
Next, we propose the quantitative requirements for the unified yet distinct latent space. To prevent features with large-scale values from overwhelming those with small-scale values, the feature densities across modalities should be similar, as formulated in (4a). Given comparable densities, the information content of each modality can be directly reflected by the feature dimensions according to (2), enabling flexible adjustment of modality contributions by modulating corresponding feature dimensions. Ideally, modality contributions in fusion should align with their confidence levels. Let $L_{img}$ and $L_{tab}$ denote the dimension of $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$, respectively, this requirement is formulated in (4b).
$\textbf{Density}(\boldsymbol{\tilde{z}}_{img})$ $\approx\textbf{Density}(\boldsymbol{\tilde{z}}_{tab})$ (4a)
$\frac{\textbf{Confidence}(\boldsymbol{\tilde{z}}_{img})}{\textbf{Confidence}(\boldsymbol{\tilde{z}}_{tab})}$ $\approx\frac{L_{img}}{L_{tab}}\approx R_{\text{conf}}$ (4b)
According to the objective outlined in (4), we propose the AMF module to modulate the original unimodal features, which constrains similar feature densities across modalities while adjusting the feature dimensions based on their confidence levels. It employs a straightforward architecture and simple training tasks to achieve accurate and efficient fusion.
@@ 3.3.2 Implementation
The structure of the AMF module is shown in the Fig. 3. First, two projector heads, $\boldsymbol{\Psi}_{img}$ and $\boldsymbol{\Psi}_{tab}$, which consistent with those used in contrastive pretraining, map the original features $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$ to the transitional features $\boldsymbol{z}\textquoteright_{img}\in\mathbb{R}^{D_{out}}$ and $\boldsymbol{z}\textquoteright_{tab}\in\mathbb{R}^{D_{out}}$. Here, $D_{out}$ denotes the input dimension of the final model. Then, a set of feature masks, leakage loss, and density loss are proposed to achieve the objective defined in (4), resulting in the flexible-dimension modulated features $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$. Finally, $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$ are concatenated to obtain the fused feature $\boldsymbol{z}_{fused}$.
Figure 3: The structure of the AMF module. For each modality, the original features are mapped to the transitional features separately. Then feature masks implicitly partition them into retained parts (solid lines) and discarded parts (dashed lines). During training, the leakage loss compresses effective information into the retained parts, and density loss aligns their density across modalities.
@@
One challenge in our framework is how to flexibly adjust the dimension of the modulated unimodal features to fit the fixed architecture of the final model, which receives the concatenated feature as input and expects a specific dimension $D_{out}$. That is, the dimensions of the modulated features, $L_{img}$ and $L_{tab}$, are constrained by $D_{out}$ as follows:
$L_{img}+L_{tab}=D_{out}.$ (5)
However, $L_{img}$ and $L_{tab}$ should vary with the $R_{conf}$ to regulate the modality contribution according to (4b). By combining (4b) and (5), specific values of $L_{img}$ and $L_{tab}$ can be derived as follows:
$\begin{cases}L_{img}&=D_{out}-\left\lfloor\frac{D_{out}}{1+R_{conf}}\right\rfloor\\ L_{tab}&=\left\lfloor\frac{D_{out}}{1+R_{conf}}\right\rfloor\end{cases}.$ (6)
Therefore, we defined a feature mask for each modality. For the imaging modality, the mask is a binary vector of dimension $D_{out}$, with the first $L_{img}$ elements set to $1$ and the rest set to 00; Conversely, the tabular feature mask is defined in the opposite way as follows:
$\begin{cases}\boldsymbol{m}_{img}&=[\underbrace{1,1,\dots,1}_{L_{img}},\underbrace{0,0,\dots,0}_{L_{tab}}]^{T}\\ \boldsymbol{m}_{tab}&=[\underbrace{0,0,\dots,0}_{L_{img}},\underbrace{1,1,\dots,1}_{L_{tab}}]^{T}\end{cases}.$ (7)
These masks divide the transitional features into a retained part and a discarded part. The retained parts are concatenated to form the fused feature, while the discarded parts which contain little information are dropped as follows:
$\boldsymbol{z}_{fused}={\boldsymbol{m}_{img}}^{T}\boldsymbol{z}\textquoteright_{img}+{\boldsymbol{m}_{tab}}^{T}\boldsymbol{z}\textquoteright_{tab}.$ (8)
@@
To prevent the discarded part from carrying away crucial information, we introduce a leakage loss that quantifies the information density in the masked-out region, as defined in (9), where $\mathcal{N}$ is the batch size.
$\mathcal{L}_{\text{leakage}}=\frac{1}{2\mathcal{N}}\sum_{j\in\mathcal{N}}\left(\frac{{\left\|{\boldsymbol{m}_{tab}}^{T}{\boldsymbol{z}\textquoteright_{img}}^{j}\right\|}_{1}}{D_{tab}}+\frac{{\left\|{\boldsymbol{m}_{img}}^{T}{\boldsymbol{z}\textquoteright_{tab}}^{j}\right\|}_{1}}{D_{img}}\right)$ (9)
By minimizing the leakage and the cross-entropy (CE) losses, both $\boldsymbol{\Psi}_{img}$ and $\boldsymbol{\Psi}_{tab}$ are trained to compress effective information into the retained part, preventing the adjustment of feature dimension from reducing feature effectiveness.
Density loss.
To align the density of the retained features as per (4a), we introduce a density loss, defined as the density difference between unimodal features in (10). By minimizing the density loss, we encourage the modulated features to exhibit similar density.
$\mathcal{L}_{\text{density}}=\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}\left|\frac{{\left\|{\boldsymbol{m}_{img}}^{T}{\boldsymbol{z}\textquoteright_{img}}^{j}\right\|}_{1}}{D_{img}}-\frac{{\left\|{\boldsymbol{m}_{tab}}^{T}{\boldsymbol{z}\textquoteright_{tab}}^{j}\right\|}_{1}}{D_{tab}}\right|$ (10)
@@ 3.4 Optimization
In the pretraining stage, image-tabular pairs are used for contrastive learning and the ITC loss is applied (Hager et al., 2023). The image-to-tabular contrastive loss is defined in (11) and the tabular-to-image contrastive loss shares a symmetric formulation.
$\mathcal{L}_{\text{i2t}}=-\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}log\frac{\text{exp}(\text{cos}({\boldsymbol{\hat{z}}_{img}}^{j},{\boldsymbol{\hat{z}}_{tab}}^{j})/\tau)}{\sum\limits_{k\in\mathcal{N},k\neq j}\text{exp}(\text{cos}({\boldsymbol{\hat{z}}_{img}}^{j},{\boldsymbol{\hat{z}}_{tab}}^{k})/\tau)}\$ (11)
Here, $\mathcal{N}$ represents the batch size, and $\tau$ is the temperature hyperparameter. The overall ITC loss is computed as the average of these two components as: $\mathcal{L}_{\text{ITC}}=\left(\mathcal{L}_{\text{i2t}}+\mathcal{L}_{\text{t2i}}\right)/2.$
In the fine-tuning stage, the loss function consists of three components. Alongside the leakage loss in (9) and density loss in (10) for multimodal fusion, we employ CE loss for classification. For the OL3I dataset, which suffers from severe class imbalance, we apply a weighted CE loss as defined in (12), where $p$ is the model output, $y$ is the classification label, and $\alpha$ is a hyperparameter for mitigate class imbalance.
@@
The overall loss function is a weighted sum of these three losses, with $\lambda_{1},\lambda_{2}$ as hyperparameters to balance leakage and density loss as follows:
$\mathcal{L}_{\text{OL3I}}$ $=\mathcal{L}_{\text{weighted ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{density}}$ (14a)
$\mathcal{L}_{\text{DVM}}$ $=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{density}}.$ (14b)
@@ 4.1 Datasets
We conducted experiments on two multimodal datasets: OL3I (Opportunistic L3 computed tomography slices for Ischemic Heart Disease risk assessment) (AIMI, 2023) and DVM (Data Visual Marketing) (Huang et al., 2021). OL3I is a medical dataset that combines grayscale computed tomography (CT) images with high-dimensional clinical tabular features, while DVM is a large-scale natural dataset for automotive applications, offering a complementary benchmark for general-purpose scenarios.
@@
The OL3I dataset contains 8,139 axial CT slices at the third lumbar vertebra (L3) level, each paired with structured medical record data. Every image is a 512$\times$512 grayscale slice, and the tabular modality includes 423 features covering patient demographics, laboratory results, and aggregated clinical codes. In our study, we adopted the 1-year IHD prediction task as the downstream objective. Due to the severe class imbalance, we used the Area Under the receiver operating Characteristic Curve (AUC) as the evaluation metric. All experiments followed the dataset’s predefined training, validation, and test splits.

### v2

@@ 3.2 FT-Mamba
Inspired by FT-Transformer [25], we adapt the Mamba architecture [8] for tabular feature extraction and name it as FT-Mamba (Feature Tokenizer + Mamba). As shown in Fig. 1 (a), we first separately tokenize the categorical and numerical tabular features to embeddings and add a [CLS] token at the sequence end. Then a stack of mamba blocks is used to extract the high-dimensional features from the embeddings, and the final representation of the [CLS] token is used as the original tabular embedding, $\boldsymbol{z}_{tab}$.
@@ 3.2.1 Feature Tokenizer
Our feature tokenizer module follows the design of [25]. It transfers the tabular input $\boldsymbol{x}_{tab}$ to embeddings $\boldsymbol{t}_{tab}\in\mathbb{R}^{N\times D_{tab}}$ as shown in Fig. 2 (a). For continuous features, the feature value is mapped by a linear function with $D_{tab}$-dimensional weights and bias. For categorical features, each category is mapped to a vector through a lookup table and is added with the corresponding bias.
@@ 3.3.1 Problem Formulation
Modality feature heterogeneity and imbalanced modality contributions can undermine the effectiveness of multimodal fusion. To address this, feature modulation aims to make the original unimodal features more compatible for fusion by mitigating heterogeneity in feature magnitude and dimensionality across modalities, while leveraging prior knowledge regarding modality confidence to adjust their contributions.
We first employ the $\ell_{1}$-norm to quantify the contribution of a unimodal feature in naive multimodal concatenation. The $\ell_{1}$-norm of a feature element directly reflects its magnitude, and the sum over all elements provides a simple and interpretable measure of the overall contribution $\mathcal{C}$ of a unimodal feature, as in (1a). Intuitively, modalities with larger overall contributions dominate in concatenated representations.
@@
To eliminate the influence of dimensionality, we define the average contribution $\bar{\mathcal{C}}$ as in (1b), which reflects the average magnitude per feature element. Equations (1a) and (1b) use $\boldsymbol{z}_{img}$ as an example, and the same concept applies to the tabular modality.
To further incorporate prior knowledge regarding modality confidence into fusion, we introduce the modality confidence ratio $r_{\text{conf}}$. This ratio represents the confidence level of image features relative to tabular features or a given downstream task, and the confidence levels can be derived from prior knowledge, such as preliminary experiments or expert experience. In our experiments, $r_{\text{conf}}$ is calculated based on unimodal performance metrics obtained from pretrained encoders on downstream tasks, denoted as Metrics:
$r_{\text{conf}}=\frac{\textbf{Metrics}(\boldsymbol{z}_{img})}{\textbf{Metrics}(\boldsymbol{z}_{tab})}.$ (2)
Based on the above definitions, we propose two objectives for modulation:
  • 1.
    Magnitude alignment. To avoid the large-magnitude features overwhelming the smaller ones, the average feature magnitude across modulated unimodal features should be comparable:
    $\left|\bar{\mathcal{C}}(\boldsymbol{\tilde{z}}_{img})-\bar{\mathcal{C}}(\boldsymbol{\tilde{z}}_{tab})\right|\rightarrow 0.$ (3)
  • 2.
    Confidence and dimensionality adjustment. To ensure that more reliable modalities contribute proportionally more to downstream tasks, the overall contribution of each modality in the fused representation should be proportional to its confidence level. Given the constraint of comparable feature magnitudes as required in (3), the overall contribution of a modality is mainly determined by its feature dimensionality according to (1b). Hence, the relative contributions can be modulated by scaling the feature lengths, which can be readily implemented within neural networks. This objective can be expressed as:
    $\frac{\mathcal{C}(\boldsymbol{z}_{img})}{\mathcal{C}(\boldsymbol{z}_{tab})}=r_{\text{conf}}\Rightarrow\frac{L_{img}}{L_{tab}}=r_{\text{conf}},$ (4)
    where $L_{img}$ and $L_{tab}$ denote the dimension of $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$, respectively.
To achieve the two objectives in (3) and (4), we propose the AMF module, which modulates the original unimodal features using a straightforward architecture and simple training tasks, enabling effective and interpretable multimodal fusion.
@@ 3.3.2 Implementation
The structure of the AMF module is shown in the Fig. 3. First, two projector heads, $\boldsymbol{\Psi}_{img}$ and $\boldsymbol{\Psi}_{tab}$, map the original features $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$ to the transitional features $\boldsymbol{z}^{\prime}_{img},\boldsymbol{z}^{\prime}_{tab}\in\mathbb{R}^{D_{out}}$, where $D_{out}$ corresponds to the dimension of the fused feature. Then, the adjustment of feature dimensions according to $r_{\text{conf}}$ is not realized by changing the structure of the projection heads. Instead, it is realized by learning appropriate parameter values during training, guided by the subsequent feature mask and leakage loss. This design enhances the stability and generalizability of the module: when $r_{\text{conf}}$ varies due to changes in downstream data quality, there is no need to modify the projection head structure, as the AMF module can adaptively modulate the features. Finally, the modulated features, $\boldsymbol{\tilde{z}}_{img}\in\mathbb{R}^{L_{img}}$ and $\boldsymbol{\tilde{z}}_{tab}\in\mathbb{R}^{L_{tab}}$, are concatenated to obtain the fused feature $\boldsymbol{z}_{fused}$ with a dimension of $D_{out}=L_{img}+L_{tab}$, which is subsequently fed into the decoder for downstream tasks.
Figure 3: The structure of the AMF module. For each modality, the original features are mapped to the transitional features separately. Then feature masks implicitly partition them into retained region (solid lines) and discarded region (dashed lines). During training, the leakage loss compresses effective information into the retained region, and magnitude loss aligns their magnitude across modalities.
@@
In practical implementation, we fix $D_{out}$ as a hyperparameter to ensure a consistent interface to downstream decoders. This design decouples the adjustment of modality contributions from the architecture of the fusion and decoding modules, thereby improving the generality of the AMF module and avoiding repeated reconfiguration of decoder input layers. Given a fixed $D_{out}$ and combined with (4), the allocation of dimensions between modalities is adaptively determined by the modality confidence ratio $r_{\text{conf}}$ as follows:
$D_{out}\;\text{is fixed},\quad\begin{cases}L_{img}(r_{\text{conf}})&=D_{out}-\left\lfloor\frac{D_{out}}{1+r_{\text{conf}}}\right\rfloor,\\ L_{tab}(r_{\text{conf}})&=\left\lfloor\frac{D_{out}}{1+r_{\text{conf}}}\right\rfloor.\end{cases}$ (5)
To explicitly adjust the dimensionalities of unimodal features, we define a feature mask for each modality to truncate the transitional features. For the imaging modality, the mask is a binary vector of size $D_{out}$, where the first $L_{img}$ entries are set to $1$ and the remaining $L_{tab}$ entries are set to 0. Conversely, the tabular feature mask adopts the opposite configuration:
$\begin{cases}\boldsymbol{m}_{img}&=[\underbrace{1,1,\dots,1}_{L_{img}},\underbrace{0,0,\dots,0}_{L_{tab}}]\\ \boldsymbol{m}_{tab}&=[\underbrace{0,0,\dots,0}_{L_{img}},\underbrace{1,1,\dots,1}_{L_{tab}}].\end{cases}$ (6)
By applying the feature masks to the transitional features, the features are divided into a retained region, $\tilde{\boldsymbol{z}}_{*}$, and a discarded region, $\tilde{\boldsymbol{z}}_{*}^{c}$, as formulated in (7). Consequently, the retained region of each modality satisfies the dimensionality objective specified in (4).
$\tilde{\boldsymbol{z}}_{img}$ $=\left[\boldsymbol{m}_{img}\odot\boldsymbol{z}^{\prime}_{img}\right]_{0:L_{img}},$ $\quad\tilde{\boldsymbol{z}}^{c}_{img}$ $=\left[\boldsymbol{m}_{tab}\odot\boldsymbol{z}^{\prime}_{img}\right]_{L_{img}:D_{out}};$ (7)
$\tilde{\boldsymbol{z}}_{tab}$ $=\left[\boldsymbol{m}_{tab}\odot\boldsymbol{z}^{\prime}_{tab}\right]_{L_{img}:D_{out}},$ $\quad\tilde{\boldsymbol{z}}^{c}_{img}$ $=\left[\boldsymbol{m}_{img}\odot\boldsymbol{z}^{\prime}_{tab}\right]_{0:L_{img}}$
Here, $\odot$ denotes the element-wise product between two vectors. During training, the projections are optimized under the guidance of the leakage and magnitude losses, which encourage modality-specific information to be concentrated in the retained region while maintaining comparable feature magnitudes. The retained regions are then used as the modulated features, which are concatenated to form the final fused representation as follows:
$\boldsymbol{z}_{fused}={\boldsymbol{m}_{img}}\odot\boldsymbol{z}^{\prime}_{img}+{\boldsymbol{m}_{tab}}\odot\boldsymbol{z}^{\prime}_{tab}.$ (8)
@@
A potential issue with masking is that meaningful information may “leak” into the discarded region, thereby reducing the effectiveness of the retained features. To address this, we introduce a leakage loss to measure the residual information magnitude within the masked-out regions:
$\mathcal{L}_{\mathrm{leakage}}=\frac{1}{2\mathcal{N}}\sum_{j\in\mathcal{N}}\left(\bar{\mathcal{C}}\big{(}\tilde{\boldsymbol{z}}_{img,j}^{c}\big{)}+\bar{\mathcal{C}}\big{(}\tilde{\boldsymbol{z}}_{tab,j}^{c}\big{)}\right)=\frac{1}{2\mathcal{N}}\sum_{j\in\mathcal{N}}\left(\frac{\left\|\left[\boldsymbol{z}^{\prime}_{img,j}\right]_{L_{img}:D_{out}}\right\|_{1}}{L_{tab}}+\frac{\left\|\left[\boldsymbol{z}^{\prime}_{img,j}\right]_{0:L_{img}}\right\|_{1}}{L_{img}}\right).$ (9)
where $\mathcal{N}$ is the batch size. By minimizing this loss jointly with the cross-entropy (CE) loss, the projection heads $\boldsymbol{\Psi}_{img}$ and $\boldsymbol{\Psi}_{tab}$ are encouraged to compress discriminative information into the retained region, ensuring that truncation by feature masks does not compromise the expressiveness of the fused features.
Magnitude loss.
To satisfy the objectives in (3), we define a magnitude loss that measures the discrepancy in information magnitude between the retained feature regions of the two modalities:
$\mathcal{L}_{\text{mag}}=\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}\left(\bar{\mathcal{C}}(\tilde{\boldsymbol{z}}_{img,j})-\bar{\mathcal{C}}(\tilde{\boldsymbol{z}}_{tab,j})\right)=\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}\left|\frac{{\left\|\left[{\boldsymbol{z}^{\prime}_{img,j}}\right]_{0:L_{img}}\right\|}_{1}}{L_{img}}-\frac{{\left\|\left[{\boldsymbol{z}^{\prime}_{tab,j}}\right]_{L_{img}:D_{out}}\right\|}_{1}}{L_{tab}}\right|.$ (10)
By minimizing this loss, the module is guided to reduce the magnitude gap across modalities, while ensuring the feasibility of adjusting the overall modality contribution through dimensional modulation.
@@ 3.4 Optimization
In the pretraining stage, image–tabular pairs are used for contrastive learning and the ITC loss is applied [6, 11]. Intuitively, the goal of ITC is to align paired image and tabular representations while pushing apart unpaired ones. The image feature of a given sample is encouraged to be close to its corresponding tabular feature, and simultaneously distant from tabular features of other samples in the batch. Formally, the image-to-tabular contrastive loss is defined in (11), where $\mathcal{N}$ is the batch size and $\tau$ is the temperature hyperparameter:
$\mathcal{L}_{\text{i2t}}=-\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}log\frac{\text{exp}(\text{cos}({\boldsymbol{\hat{z}}_{img}}^{j},{\boldsymbol{\hat{z}}_{tab}}^{j})/\tau)}{\sum\limits_{k\in\mathcal{N},k\neq j}\text{exp}(\text{cos}({\boldsymbol{\hat{z}}_{img}}^{j},{\boldsymbol{\hat{z}}_{tab}}^{k})/\tau)}\\$ (11)
The tabular-to-image contrastive loss adopts a symmetric formulation, and the overall ITC loss is computed as their average as: $\mathcal{L}_{\text{ITC}}=\left(\mathcal{L}_{\text{i2t}}+\mathcal{L}_{\text{t2i}}\right)/2.$
In the fine-tuning stage, the loss function consists of three components. Alongside the leakage loss in (9) and magnitude loss in (10) for multimodal fusion, we employ CE loss for classification. For the OL3I dataset, which suffers from severe class imbalance, we apply a weighted CE loss as defined in (12), where $p$ is the model output, $y$ is the classification label, and $\alpha$ is a hyperparameter for mitigate class imbalance.
@@
The overall loss function is a weighted sum of these three losses, with $\lambda_{1},\lambda_{2}$ as hyperparameters to balance leakage and magnitude loss as follows:
$\mathcal{L}_{\text{OL3I}}$ $=\mathcal{L}_{\text{weighted ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{magnitude}}$ (14a)
$\mathcal{L}_{\text{DVM}}$ $=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{magnitude}}.$ (14b)
@@ 4.1 Datasets
We conducted experiments on two multimodal datasets: OL3I (Opportunistic L3 computed tomography slices for Ischemic Heart Disease risk assessment) [29] and DVM (Data Visual Marketing) [30]. OL3I is a medical dataset that combines grayscale computed tomography (CT) images with high-dimensional clinical tabular features, while DVM is a large-scale natural dataset for automotive applications, offering a complementary benchmark for general-purpose scenarios.
@@
The OL3I dataset contains 8,139 axial CT slices at the third lumbar vertebra (L3) level, each paired with structured medical record data. Every image is a $\times512$ grayscale slice, and the tabular modality includes 423 features covering patient demographics, laboratory results, and aggregated clinical codes. In our study, we adopted the 1-year IHD prediction task as the downstream objective. Due to the severe class imbalance, we used the Area Under the receiver operating Characteristic Curve (AUC) as the evaluation metric. All experiments followed the dataset’s predefined training, validation, and test splits.